# Unveiling the Influence of Amplifying Language-Specific Neurons

### v2

@@ Abstract
Language-specific neurons in LLMs that strongly correlate with individual languages have been shown to influence model behavior by deactivating them. However, their role in amplification remains underexplored. This work investigates the effect of amplifying language-specific neurons through interventions across 18 languages, including low-resource ones, using three models primarily trained in different languages. We compare amplification factors by their effectiveness in steering to the target language using a proposed Language Steering Shift (LSS) evaluation score, then evaluate it on downstream tasks: commonsense reasoning (XCOPA, XWinograd), knowledge (Include), and translation (FLORES). The optimal amplification factors effectively steer output toward nearly all tested languages. Intervention using this factor on downstream tasks improves self-language performance in some cases but generally degrades cross-language results. These findings highlight the effect of language-specific neurons in multilingual behavior, where amplification can be beneficial especially for low-resource languages, but provides limited advantage for cross-lingual transfer.  [^1] The code and dataset are made available at https://github.com/tauimbz/lang-task-neuron.
@@ 1 Introduction
In elucidating that issue, a few recent studies have begun to interpret model inner workings in multilingual contexts, revealing neuron-level correlations with language features and suggesting the existence of language-specific neurons Kojima et al. (2024); Tang et al. (2024); Zhao et al. (2024). These neurons have been shown to affect multilingual capability following deactivation and signs of language-steering capabilities in generation. Interventions typically use fixed activation values, called steering factor, to control behavior. Most prior works use zero as the steering factor to deactivate targeted neurons. However, the impacts of actively steering these neurons in downstream tasks, particularly related to improving performance or enabling task transfer, remain unclear as recent studies have focused only on self-language interventions Mondal et al. (2025); Wang et al. (2024).
This work addresses the gap by identifying and analyzing two types of neurons: language-specific neurons and language-activated (less-specific) neurons. We then compare various steering factors derived from different settings to evaluate their effect on the steerability of these neurons and analyze their impact on downstream task performance using neurons specific to the target language (self-intervention) and to other languages (cross-intervention). We evaluate 18 languages including low-resource ones like Quechua, Haitian Creole, and Swahili, across models primarily trained in English and Chinese (Qwen), English (Gemma), and Southeast Asian languages (SeaLLMs) using XCOPA, XWinograd, and Include-lite for reasoning and knowledge tasks, and FLORES-200 for translation.
@@
  • Comparison of different steering factors using the proposed Language Steering Shift (LSS) reveals that patched steering factors are the most effective and statistically outperform test-time intervention factors. Amplifying language-specific neurons with these factors consistently directs outputs toward all target languages, achieving over 90% average success in language shifting.
[…]
@@ 3.1 Neuron Identification
language activation probablity entropy (LAPE) method (Tang et al., 2024) was adopted to identify language-specific neurons. The probability of activation for the $j$-th neuron within the $i$-th layer when processing language $k$, denoted by $p^{k}_{i,j}$, is estimated according to:
@@
The expectation is taken over all tokens from sentences in language $k$, where $\mathbb{I}(.)$ is an indicator function that returns 111 if the condition is true and 0 otherwise; and $\texttt{act}(i,j)$ denotes the activation value of the $j$-th neuron in the $i$-th layer. Let $\mathbf{p}_{i,j}=[p^{1}_{i,j},\dots,p^{k}_{i,j},\dots,p^{|L|}_{i,j}]$ represent a list of activation probabilities for a set of languages $L$. These probabilities are then normalized to yield the L1-norm version, denoted by $p^{\prime}_{i,j}$. Using this normalized version, the LAPE score of the $j$-th neuron in the $i$-th layer is formally defined as:
@@ 3.3 Determining steering factors
Patched Steering Factors. Let $v_{j,k}$ be the variable that contains the activation value of neuron $j$ that is specific to language $k$. In the identification stage involving the multilingual FLORES-200 dataset, the steering factors are computed from the mean and median of activation values across all sentences within the dataset $S_{k}$ for language $k$. As a result, a steering factor is specific to neuron $j$ and language $k$, to which we patch the language-specific neurons. This process is formally described as:
@@
[…]
@@
[…]
@@
3.4 LSS Scores and Optimal Amplifying Steering Factor
Result of the Language Steering Shift (LSS) score shows that LAPE neuron amplification yields high scores for many languages, often reaching 100%, as shown in Table 1 for the intervention on id. These strong results are consistent across other languages, as detailed in Appendix E. However, interventions involving certain language pairs may result in lower LSS scores. For example, translation between id and ms in either direction tends to produce scores capped at 50% in most models. This is likely due to the high similarity between the two languages, which makes it more difficult for the models to distinguish between them. Notably, en remains easy to steer to, suggesting a strong influence that is possibly due to its dominance and widespread presence in LLM training data.
[…]
@@ 4 Experimental Settings
We identify language-activated (Baseline) and language-specific (LAPE) neurons by feeding the LLMs sentences from the dev split of FLORES-200 dataset Team et al. (2022) containing 30,00030{,}00030 , 000 word-tokens per language. We assess 18 languages to ensure broader coverage of language specificity: en, nl, id, ms, vi, jp, zh, fr, pt, ru, et, ht, it, qu, sw, ta, th, and tr in small and large models, namely Qwen2.5 (0.5B and 7B), Gemma2 (2B and 9B), and SeaLLMs-v3 (1.5B and 7B), totaling 6 models (see Appendix B for details). For identifying LAPE neurons, we use a language activation probability threshold ($m$) set at the 95th percentile and a LAPE score threshold ($n$) of 1%. All hyperparameter values are adopted from Tang et al. (2024) as they have been demonstrated to effectively capture language-specific neurons.
@@
We assess the determined steering factors on LSS score for the small models using a reconstructed MLAMA Kassner et al. (2021), a multilingual fact-probing dataset, so that each language data includes the answers of every other languages and filters the rows to ensure that no answers in different languages were the same. 15 out of 18 languages are used due to limitation in the dataset.
@@
We evaluate the steering factor for commonsense reasoning, which is considered one of the most difficult tasks for LLMs Li et al. (2022), using XCOPA Edoardo M. Ponti and Korhonen (2020) and XWinograd Muennighoff et al. (2022). For factual knowledge, we use Include-lite Romanou et al. (2024). Building on Mondal et al. (2025), who focused only on self-language intervention, we extend the analysis to cross-lingual intervention.
The devtest split of FLORES-200 is used to measure perplexity and BLEU for translation from English (en) to every other languages. Translation tasks are performed using 2 types of prompts: targeted and non-targeted. Targeted prompts explicitly mention the target language, whereas non-targeted prompts do not. For example, when translating to id, the targeted prompt begins with "Translate from English to Indonesian.", whereas the non-targeted prompt starts with "Translate from English into the target language."
@@ 5 Result and Analysis
This section presents the findings and analysis from our steering factor evaluations.
@@ D O D O D O
S-7B $\mathbf{0.38}$ 0.11-0.11- 0.11 $\mathbf{\phantom{-}1.27}$ 6.71-6.71- 6.71 $\mathbf{-0.44}$ 5.49-5.49- 5.49
[…]
Q-7B $\mathbf{0.26}$ 0.17-0.17- 0.17 $\mathbf{-0.03}$ 7.19-7.19- 7.19 $\mathbf{-0.22}$ 5.98-5.98- 5.98
Q-0.5B $\mathbf{0.16}$ 0.36-0.36- 0.36 $\mathbf{\phantom{-}0.17}$ 3.38-3.38- 3.38 $\mathbf{-0.04}$ 1.69-1.69- 1.69
G-9B $\mathbf{0.87}$ 0.77-0.77- 0.77 $\mathbf{\phantom{-}0.29}$ 5.46-5.46- 5.46 $\mathbf{\phantom{-}0.13}$ 2.24-2.24- 2.24
G-2B $\mathbf{0.25}$ 1.03-1.03- 1.03 $\mathbf{\phantom{-}0.15}$ 3.38-3.38- 3.38 $\mathbf{\phantom{-}0.07}$ 1.43-1.43- 1.43
Table 2: Average delta tasks accuracy after pmax intervention on LAPE neurons for self-intervention/diagonal (D) and cross-intervention/off-diagonal (O) neuron steering across models. S denotes SeaLLMs3, Q = Qwen2.5 and G = Gemma2.
@@ 5.1 Impact of Neuron Amplification
Figure 3: Delta Include-lite accuracy under LAPE neurons on Gemma2 9B.
@@ 5.1.1 Knowledge and Reasoning Tasks
Overall performance of knowledge and reasoning tasks intervention as presented in Table 2 shows that pmax self-intervention can improve accuracy up to 10%, with consistent gains observed in Include-lite. XCOPA and XWinograd self-intervention may show less consistent improvements with only modest changes in some models. In contrast, cross-language interventions generally lead to performance drops on average. These results suggest that amplifying LAPE neurons in self-language on knowledge and reasoning tasks is potentially beneficial for enhancing performance, though their effectiveness in cross-lingual transfer appears limited.
@@
Figure 4 shows that self-intervention tends to be non-destructive, with some languages such as jp and pt, showing slight improvements across most models. However, the improved languages vary by models; for example, zh improves in Qwen2.5-0.5B but not in the 7B variant. Cross-lingual interventions degrade performance on average, particularly in larger models. Nevertheless, occasional improvements are observed in less related language pairs, such as zh when intervened by ru in Qwen2.5-0.5B (see Appendix H for additional results). A consistent trend across models is that amplifying en neurons reliably reduces accuracy, suggesting they may play a role in reasoning rather than language transfer.
@@
Figure 5 shows that amplifying LAPE neurons does not consistently improve in self-intervention, but changes are generally reserved, indicating non-destructive effects. In contrast, cross-language amplification often leads to larger drops, visibly captured on the intervention of it, id, th, tr, and vi neurons, degrading other languages more than their own. Low-resource languages like qu, et, sw, and ta show self-intervention gains, particularly in SeaLLMs3. SeaLLMs3 and Qwen2.5, having shared architecture Zhang et al. (2024), display very similar patterns, suggesting architecture may contribute to neuron behavior. Overall, XCOPA deltas reveal clearer signs of degradation than improvement.
@@ 5.1.2 Translation Task
Figure 3 shows that BLEU scores increase after amplifying LAPE neurons with pmax for translating English to 17 other languages with non-targeted prompt. The results suggest that amplifying language-specific neurons enhances the model’s ability to generate text in the target language, including low-resource languages such as qu, ht, and sw, across most models. In contrast, the improvements from targeted prompts are less pronounced. Smaller models like Qwen2.5 0.5B and SeaLLMs3-1.5B tend to exhibit more consistent gains, which may indicate a higher degree of neuron locality in these models. The difference in performance between targeted and non-targeted prompts may suggest that amplifying these neurons improves the generation of the target language, though not always with high accuracy for translation. In contrast to amplification of Baseline neurons, which often degrades BLEU scores (Appendix G), LAPE demonstrates the ability to improve performance by targeting language-specific neurons.
Figure 6: BLEU deltas for non-targeted pmax intervention in Qwen2.5 0.5B. Row $i$ is the target language and column $j$ is the intervention language. For example, row $x$ col $y$ represents BLEU score when translating English to language $x$ intervened by $y$ neurons.
To support the increasing BLEU findings, particularly in non-targeted prompt, we compare the performance with the cross-intervention setting presented in Figure 6. Self-intervention results in increasing BLEU scores compared to cross-intervention. Notably, using ru as the intervention language reduces performance in the cross-intervention case, while other languages show more modest increase compared to self-intervention. The observed improvements across languages may relate to linguistic similarity. For example, some increases occur when translating to id with ms as the intervention, and vice versa; or when translating to pt with fr or it, and to fr with ht.

### v3

@@ Abstract
Language-specific neurons, units in LLMs that strongly correlate with individual languages have been shown to influence model behavior by deactivating them. However, their role in amplification remains underexplored. This work investigates the effect of amplifying language-specific neurons through interventions across 18 languages, including low-resource ones, using three models primarily trained in different languages. We compare amplification factors by their effectiveness in steering to the target language using a proposed Language Steering Shift (LSS) evaluation score, then evaluate it on downstream tasks: commonsense reasoning (XCOPA, XWinograd), knowledge (Include), and translation (FLORES). The optimal amplification steering factors effectively steer output toward nearly all tested languages. Intervention using this factor on downstream tasks improves self-language performance in some cases but generally degrades cross-language results. These findings highlight the effect of language-specific neurons in multilingual behavior, where amplification can be beneficial especially for low-resource languages, but provides limited advantage for cross-lingual transfer.  [^1] The code and dataset are made available at https://github.com/tauimbz/lang-task-neuron.
@@ 1 Introduction
In elucidating this issue, a few recent studies have begun to interpret model inner workings in multilingual contexts, revealing neuron-level correlations with language features and suggesting the existence of language-specific neurons Kojima et al. (2024); Tang et al. (2024); Zhao et al. (2024). These neurons have been shown to affect multilingual capability following deactivation and signs of language-steering capabilities in generation. Interventions are typically done to steer outputs by replacing the neuron values with fixed activation values, called steering factor, to control behavior. Most prior works use zero as the steering factor to deactivate targeted neurons. However, we argue that amplifying these neurons can offer deeper insight into their functional capabilities, particularly regarding performance improvements or task and language transfer.
Building on the success of cross-lingual transfer methods such as training and fine tuning Conneau et al. (2020); Kim et al. (2019), we aim to investigate whether neuron overlap reflects shared linguistic representations. To this end, we extend existing studies that have been largely focusing on self-language interventions Mondal et al. (2025); Wang et al. (2024) which showed limited gains, with cross-language interventions. We examine cross-lingual neuron behavior in closely related language pairs, distant language pairs, and low resource languages, to further explore whether neuron-level interventions can similarly provide insights on cross-lingual transfer, particularly revealing where and why such interventions succeed or fail.
This work is conducted by identifying and analyzing two types of neurons: language-specific neurons and language-activated (less-specific) neurons. We then compare various steering factors derived from different settings to evaluate their effect on the steerability of these neurons and analyze their impact on downstream task performance using neurons specific to the target language (self-intervention) and to every other languages (cross-intervention). We evaluate 18 languages including low-resource ones like Quechua, Haitian Creole, and Swahili, across models primarily trained in English and Chinese (Qwen), English (Gemma), and Southeast Asian languages (SeaLLMs) using XCOPA, XWinograd, and Include-lite for reasoning and knowledge tasks, and FLORES-200 for translation.
@@
  • Comparison between different steering factors using the proposed Language Steering Shift (LSS) reveals that patched steering factors are the most effective and statistically outperform test-time intervention factors. Amplifying language-specific neurons with these factors consistently directs outputs toward all target languages, achieving over 90% average success in language shifting.
  • Optimal amplification of language-specific neurons in self-intervention can reduce perplexity and improve task performance for certain languages, with gains reaching up to $1\%$ in some models, showing improved results for low-resource languages compared to Mondal et al. (2025).
@@ 3.1 Neuron Identification
language activation probablity entropy (LAPE) method (Tang et al., 2024) is adopted to identify language-specific neurons. The probability of activation for the $j$-th neuron within the $i$-th layer when processing language $k$, denoted by $p^{k}_{i,j}$, is estimated according to:
@@
The expectation is taken over all tokens from sentences in language $k$, where $\mathbb{I}(.)$ is an indicator function that returns $1$ if the condition is true and 0 otherwise; and $\texttt{act}(i,j)$ denotes the activation value of the $j$-th neuron in the $i$-th layer. Let $\mathbf{p}_{i,j}=[p^{1}_{i,j},\dots,p^{k}_{i,j},\dots,p^{|L|}_{i,j}$] represent a list of activation probabilities for a set of languages $L$. These probabilities are then normalized to yield the L1-norm version, denoted by $p^{\prime}_{i,j}$. Using this normalized version, the LAPE score of the $j$-th neuron in the $i$-th layer is formally defined as:
@@ 3.3 Determining steering factors
Patched Steering Factors. Let $v_{j,k}$ be the variable that contains the activation value of neuron $j$ that is specific to language $k$. In the identification stage involving the multilingual FLORES-200 dataset, the steering factors are computed from the max and median of activation values across all sentences within the dataset $S_{k}$ for language $k$. As a result, a steering factor is specific to neuron $j$ and language $k$, to which we patch the language-specific neurons. This process is formally described as:
@@
where $\texttt{agg}\in\{\texttt{max},\texttt{median}\}$; and $a_{j,s}$ represents the activation value of neuron $j$ averaged across all tokens in sentence $s$ during the identification stage. We follow Kojima et al. (2024) in using the median and extend their approach by incorporating a stronger aggregation function, max. We refer to these steering factors as patch max (pmax) and median (pmedian).
@@
where $\texttt{steer}_{j,m}$ is the steering factor for neuron $j$ during inference on sentence $m$; $T(m)$ is the set of tokens in sentence $m$; and $a_{j,m,t}$ is the activation value of neuron $j$ associated with token $t$ on sentence $m$. As a comparison to our amplifying factors, we also use deactivating factors 0 and $10p$ (10th percentile) as a replacement of max, which has been similarly demonstrated by Mondal et al. (2025) to deactivate neurons. We refer to these as =0 and =10p, respectively.
@@ 4 Experimental Settings
This section discusses the experimental settings used throughout the methodology. It consists of three core stages. First, the identification of neurons specifically activated by language and those exhibiting language-specific responses. Second, the process of finding the optimal amplifying steering factor for influencing these neural activities. Finally, the evaluation stage to determine the impact of this optimal amplifying steering factor on various downstream tasks to assess its efficacy.
@@
We identify language-activated (Baseline) and language-specific (LAPE) neurons by feeding the LLMs sentences from the dev split of FLORES-200 dataset Team et al. (2022) containing 30,00030{,}000 word-tokens per language. We assess 18 languages to ensure broader coverage of language specificity: en, nl, id, ms, vi, jp, zh, fr, pt, ru, et, ht, it, qu, sw, ta, th, and tr in small and large models, namely Qwen2.5 Instruct (0.5B and 7B), Gemma2 Instruct(2B and 9B), and SeaLLMs3 Chat (1.5B and 7B), totaling in 6 models (see Appendix B for details). For identifying LAPE neurons, we use a language activation probability threshold ($m$) set at the 95th percentile and a LAPE score threshold ($n$) of 1%. All hyperparameter values are adopted from Tang et al. (2024) as they have been demonstrated to effectively capture language-specific neurons.
@@
We assess the determined steering factors on LSS score for the small models using a reconstructed version of MLAMA Kassner et al. (2021), a multilingual fact-probing dataset. MLAMA is reconstructed such that each language question includes answers from all other languages and is filtered to ensure that no answer in different languages are the same. For example, if both Indonesian and Malaysian answers are “Jerman” (i.e., Germany in English), that row is discarded. The dataset contains 185 instances per language, with 15 of the 18 available languages are used due to limitation in the dataset.
@@
We evaluate the model’s confidence and downstream task performances by conducting interventions using the optimal amplifying steering factors. We provide the deltas between the intervened and non-intervened performances to quantify the extent to which the interventions shift task outcomes. Interventions are performed in every layer of the model, with layer-wise amplification result and analysis provided in Appendix A.1.
We evaluate on commonsense reasoning task, which is considered one of the most difficult tasks for LLMs Li et al. (2022), using XCOPA Edoardo M. Ponti and Korhonen (2020) and XWinograd Muennighoff et al. (2022) dataset. For factual knowledge, we use Include-lite Romanou et al. (2024) dataset. Building on Mondal et al. (2025), who focused only on self-language intervention, we extend the analysis of these task performances to cross-lingual intervention. The devtest split of FLORES-200 is used to measure perplexity and BLEU for translation from English (en) to every other languages.
Translation tasks are performed using 2 types of prompts: targeted and non-targeted. Targeted prompts explicitly mention the target language, whereas non-targeted prompts do not. For example, when translating to id, the targeted prompt begins with "Translate from English to Indonesian.", whereas the non-targeted prompt begins with "Translate from English into the target language."
@@ 5 Result and Analysis
This section presents the findings and analysis of the optimal amplifying steering factors and their evaluation on model confidence and downstream tasks.
5.1 LSS Scores and Optimal Amplifying Steering Factor
Result of the Language Steering Shift (LSS) score shows that LAPE neuron amplification yields high scores for many languages, often reaching 100%, as shown in Table 1 for the intervention on id. These strong results are consistent across other languages, as detailed in Appendix E, indicating successful language shifting. However, interventions involving certain language pairs may result in lower LSS scores. For example, translation between id and ms in either direction tends to produce scores capped at 50% in most models. This is likely due to the high similarity between the two languages, which makes it more difficult for the models to distinguish between them. Notably, en remains easy to steer to, suggesting a strong influence that is possibly due to its dominance and widespread presence in LLM training data.
The optimal amplifying factors, defined as those that result in the highest aggregated LSS scores, are found to be pmax and pmedian. A one-way ANOVA followed by Tukey’s HSD test for post-hoc pairwise comparisons reveals that the patched steering factors pmax (mean = 0.9204) and pmedian ($\texttt{mean}=0.9095$) significantly outperform the other factors ($p<0.001$); Further results are provided in Appendix E. However, the difference between the two is not statistically significant ($p=0.98$), indicating that both are equally effective as optimal amplifying factors.
Test-time intervention factors, +max and =max, perform well mainly for high-resource languages such as en, pt, it, and fr, but show limited effectiveness for other languages, with the lowest performance observed for ta in Qwen and Gemma despite its substantial neuron presence. The deactivating factor =0 shows variation in results, with successful steering in some languages (e.g., en, nl, and it), while =10p demonstrates stronger deactivation capability, often resulting in LSS scores close to $0\%$ across many languages, align with findings from Mondal et al. (2025).
@@ D O D O D O
S-7B $\mathbf{0.38}$ 0.11-0.11 $\mathbf{\phantom{-}1.27}$ 6.71-6.71 $\mathbf{-0.44}$ 5.49-5.49
S-1.5B $\mathbf{1.02}$ $\phantom{-}0.54$ $\mathbf{-0.76}$ 6.83-6.83 $\mathbf{-0.18}$ 3.18-3.18
Q-7B $\mathbf{0.26}$ 0.17-0.17 $\mathbf{-0.03}$ 7.19-7.19 $\mathbf{-0.22}$ 5.98-5.98
Q-0.5B $\mathbf{0.16}$ 0.36-0.36 $\mathbf{\phantom{-}0.17}$ 3.38-3.38 $\mathbf{-0.04}$ 1.69-1.69
G-9B $\mathbf{0.87}$ 0.77-0.77 $\mathbf{\phantom{-}0.29}$ 5.46-5.46 $\mathbf{\phantom{-}0.13}$ 2.24-2.24
G-2B $\mathbf{0.25}$ 1.03-1.03 $\mathbf{\phantom{-}0.15}$ 3.38-3.38 $\mathbf{\phantom{-}0.07}$ 1.43-1.43
Table 2: Average delta task accuracies after pmax intervention on LAPE neurons for self-intervention/diagonal (D) and cross-intervention/off-diagonal (O) neuron steering for every models used. Negative and positive values respectively represent decreased and increased accuracy from the baseline (non-intervention scenario). S denotes SeaLLMs3, Q = Qwen2.5 and G = Gemma2.
@@ 5.2 Impact of Neuron Amplification
Figure 3: Delta Include-lite accuracy under LAPE neurons on Gemma2 9B. Blue highlights reductions, whereas red highlights increases.
@@ 5.2.1 Knowledge and Reasoning Tasks
Overall performance of knowledge and reasoning tasks intervention as presented in Table 2 shows that pmax self-intervention can improve accuracy up to 1%, with consistent gains observed in Include-lite. XCOPA and XWinograd self-intervention may show less consistent improvements with only modest changes in some models. In contrast, cross-language interventions generally lead to performance drops on average. These results suggest that amplifying LAPE neurons in self-language on knowledge and reasoning tasks is potentially beneficial for enhancing performance, though their effectiveness in cross-lingual transfer appears limited.
@@
Figure 4 shows that self-intervention tends to be non-destructive, with some languages such as jp and pt, showing slight improvements across most models. However, the improved languages vary by models; for example, zh improves in Qwen2.5 0.5B but not in the 7B variant. Cross-lingual interventions degrade performance on average, particularly in larger models. Nevertheless, occasional improvements are observed in less related language pairs, such as zh when intervened by ru in Qwen2.5 0.5B (see Appendix H for additional results). A consistent trend across models is that amplifying en neurons reliably reduces accuracy, suggesting they may play a role in reasoning rather than language transfer.
@@
Figure 5 shows that amplifying LAPE neurons does not consistently improve in self-intervention, but changes are generally reserved, indicating non-destructive effects. In contrast, cross-language amplification often leads to larger drops, visibly captured on the intervention of it, id, th, tr, and vi neurons, degrading other languages more than their own. Low-resource languages like qu, et, sw, and ta show self-intervention gains, particularly in SeaLLMs3. SeaLLMs3 and Qwen2.5, having shared model architecture Zhang et al. (2024), display very similar patterns, suggesting architecture may contribute to neuron behavior. Overall, XCOPA deltas reveal clearer signs of degradation than improvement.
@@
Figure 6: BLEU deltas for non-targeted pmax intervention in Qwen2.5 0.5B. Row $i$ is the target language and column $j$ is the intervention language. For example, row $x$ col $y$ represents BLEU score when translating English to language $x$ intervened by $y$ neurons.Blue highlights reductions, whereas red highlights increases.
@@ 5.2.2 Translation Task
Table 3 shows that BLEU scores increase after amplifying LAPE neurons with pmax for translating English to 17 other languages with non-targeted prompt. The results suggest that amplifying language-specific neurons enhances the model’s ability to generate text in the target language, including low-resource languages such as qu, ht, and sw, across most models. In contrast, the improvements from targeted prompts are less pronounced. Smaller models like Qwen2.5 0.5B and SeaLLMs3 1.5B tend to exhibit more consistent gains, which may indicate a higher degree of neuron locality in these models. The difference in performance between targeted and non-targeted prompts may suggest that amplifying these neurons improves the generation of the target language, though not always with high accuracy for translation. In contrast to amplification of Baseline neurons, which often degrades BLEU scores (Appendix G), LAPE demonstrates the ability to improve performance by targeting language-specific neurons.
To support the increasing BLEU findings, particularly in non-targeted prompt, we compare the performance with the cross-intervention setting presented in Figure 6. Self-intervention results in increasing BLEU scores compared to cross-intervention. Notably, using ru as the intervention language reduces performance in the cross-intervention case, while other languages show more modest increase compared to self-intervention. The observed improvements across languages may relate to linguistic similarity. For example, some increases occur when translating to id with ms as the intervention language, and vice versa; or when translating to pt with fr or it, and to fr with ht.